OpenStudy (anonymous):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can make as many corrections as she wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Diana can receive. Identify this as converging or diverging.

OpenStudy (anonymous):

For an infinite geometric series, the sum is \[\sum_{n=1}^{\infty}ar ^{n-1}=\frac{a}{1-r}\] first term = a= 65, common ratio = r = 0.3 \[\sum_{n=1}^{\infty}65(0.3) ^{n-1}=\frac{65}{1-0.3}\]